OpenStudy (anonymous):

Solve the equation for 0≤x<2π » 2sin²2x=1

OpenStudy (vishal_kothari):

2sin²2x = 1 divide both sides by 2 sin²2x = 1/2 take the sq roots of both the sides sin2x = -1/√2 or sin2x = 1/√2 Take the inverse sine of both sides: 2x = -π/4 or sin2x = 1/√2 divide both sides by 2 x = -π/8 or sin2x = 1/√2 take the inverse of sine of both sides x = -π/8 or 2x = π/4 divide both sides by 2 x = -π/8 or x = π/8...